Pollsters are concerned about declining levels of cooperation among persons contacted in surveys.
A pollster contacts 76 people in the 18-21 age bracket and finds that 65 of them respond and 11 of them refuse to respond.
When 303 people in the 22-29 age bracket are contacted, 283 respond and 20 refuse to respond.
Assume that 1 of the 379 people is randomly selected.
Find the probability of getting someone in the 18-21 age bracket or someone that does not respond.
Answer to a percent rounded to 1 decimal.

Answer by math_tutor2020(3817)


Table:
ResponseNo responseTotal
Age 18 to 21651176
Age 22 to 2928320303
Total34831379


Define the two events
A = someone is in the 18-21 age bracket
B = the person does not respond

Based on the table, we have
n(A) = number of people in set A = 76
n(B) = 31
n(A and B) = 11

So,
n(A or B) = n(A) + n(B) - n(A and B)
n(A or B) = 76+31-11
n(A or B) = 96
We have 96 people in the age 18-21 bracket, they didn't respond, or both.

Divide this over the 379 people surveyed
96/379 = 0.253 = 25.3%

Answer: 25.3%